# Algebraization of Mochizuki’s anabelian variation of ring structures, perfectoid geometry and formal groups

### v4

@@ Abstract
Let $M$ be a multiplicative monoid with identity. Then I show that there is a universal one dimensional formal group law equipped with an action of $M$. If $M$ is $p$-perfect (i.e. $m\longmapsto m^{p}$ is an isomorphism for some prime number $p$) then the universal $M$-formal group law comes equipped with a natural Frobenius endomorphism. There are a number of concrete applications of this result. If $K$ is a $p$-adic field and ${\mathscr{O}}={\mathscr{O}}_{K}$ is the multiplicative monoid of the ring of integers of $K$, then there is a universal formal group (over a suitable (non-zero) ring) which is equipped with an action of the multiplicative monoid ${\mathscr{O}}$. Lubin-Tate formal groups arise from this universal monoid formal group law. This has applications to Mochizuki’s anabelian ideas: if two p-adic fields have isomorphic absolute Galois groups then they have isomorphic multiplicative monoids ${\mathscr{O}}$ (but possibly non-isomorphic ring structures). The existence of the universal monoid formal group law for the monoid ${\mathscr{O}}$ implies that the additive structures of a ring can be interpolated into a universal algebraic family (while keeping the multiplicative structure of the ring fixed). Here is another important example covered by my result: let $R$ be a perfectoid ring and let $R^{\flat}$ be its tilt and the multiplicative monoid $R^{\flat}$ of $R^{\flat}$. Then there exists a universal monoid formal group law for this monoid which interpolates the additive structures of untilts with tilt $R^{\flat}$. Thus in some sense one has a unified approach to various phenomenon which are well-known in anabelian geometry and in perfectoid geometry.
@@ 1.2 The key insight
Importantly, not only can this problem be algebraized, my approach also provides a unified view of both anabelian fluidity of ring structures and similar fluidity of ring structures (with fixed multiplicative structures) encountered in perfectoid geometry (Scholze, 2012), (Kedlaya and Liu, 2015) and (Fargues and Fontaine, 2018).
@@ 1.4 Acknowledgments
I began thinking about the problem addressed in this paper while I was enjoying the support and hospitality of Research Institute of Mathematical Sciences (RIMS), Kyoto and I thank RIMS for the same. I thank Shinichi Mochizuki for conversations around many topics of common interest during my visit. I thank Yuichiro Hoshi for conversations about Anabelian Reconstruction Theory. I thank Machiel van Frankenhuijsen for many conversations related to Mochizuki’s work and the $abc$-conjecture. I thank Taylor Dupuy and Anton Hilado for some conversations surrounding Mochizuki’s works and for pointing out some typos in an earlier version of this paper. Finally it is a pleasure to thank Michael Hopkins for his insightful lectures at the Arizona Winter School 2019 and the organizers of the School for an excellently organized meeting.
@@ 2.3 Existence of formal groups with monoid actions
Let ${\mathbb{Q}}[M]$ be the monoid ring of a monoid $M$ over ${\mathbb{Q}}$. Let $R\supset{\mathbb{Q}}[M]$ be the monoid ring over ${\mathbb{Q}}$. Then I claim that there always exists an $M$-formal group over $R$.
@@ 4.2 Perfectoid rings, fields and Huber pairs and monoid formal group laws
This demonstrates the close relationship between Mochizuki’s idea of variation of ring structures keeping multiplicative structures fixed and the ideas in perfectoid geometry of (Scholze, 2012), (Kedlaya and Liu, 2015) (the relationship with (Fargues and Fontaine, 2018) is detailed in (Joshi, 2023a) and will not be discussed here). This recognition played an important role in the theory of Arithmetic Teichmuller Spaces documented extensively in (Joshi, 2021, 2023b, 2023a, 2023c). In this section, I record these observations.

### v6

@@ Abstract
Let $M$ be a multiplicative monoid with identity. Then I show that there is a universal one dimensional formal group law equipped with an action of $M$. If $M$ is $p$-perfect (i.e. $m\longmapsto m^{p}$ is an isomorphism for some prime number $p$) then the universal $M$-formal group law comes equipped with a natural Frobenius endomorphism. There are a number of concrete applications of this result. If $K$ is a $p$-adic field and ${\mathscr{O}}={\mathscr{O}}_{K}$ is the multiplicative monoid of the ring of integers of $K$, then there is a universal formal group (over a suitable (non-zero) ring) which is equipped with an action of the multiplicative monoid ${\mathscr{O}}$. Lubin-Tate formal groups arise from this universal monoid formal group law. This has applications to Mochizuki’s anabelian ideas: if two p-adic fields have isomorphic absolute Galois groups then they have isomorphic multiplicative monoids ${\mathscr{O}}$ (but possibly non-isomorphic ring structures). The existence of the universal monoid formal group law for the monoid ${\mathscr{O}}$ implies that the additive structures of a ring can be interpolated into a universal algebraic family (while keeping the multiplicative structure of the ring fixed). Here is another important example covered by my result: let $R$ be a perfectoid ring and let $R^{\flat}$ be its tilt and the multiplicative monoid $R^{\flat}$ of $R^{\flat}$. Then there exists a universal monoid formal group law for this monoid which interpolates the additive structures of untilts with tilt $R^{\flat}$. Thus in some sense one has a unified approach to various phenomenon which are well-known in anabelian geometry and in perfectoid geometry. These results also provide a natural number field version of Fontaine’s fundamental ring $A_{inf}$ of $p$-adic Hodge Theory (Subsection 4.3).
@@ 1.2 The key insight
Importantly, not only can this problem be algebraized, my approach also provides a unified view of both anabelian fluidity of ring structures and similar fluidity of ring structures (with fixed multiplicative structures) encountered in perfectoid geometry (Scholze, 2012), (Kedlaya and Liu, 2015) and (Fargues and Fontaine, 2018), and also provides similar objects in the global arithmetic case of varieties over number fields and also provides the global ring $\mathbb{A}_{inf,L}$ which is the natural analog of Fontaine’s fundamental ring $A_{inf}$ of $p$-adic Hodge Theory ((Fontaine, 1994)).
@@ 1.3 Results proved in this paper
In subsection 4.3, using the results of Subsection 4.1 and Subsection 4.2, I sketch the construction of the global (i.e. number field) analog $\mathbb{A}_{inf,L}$ of the fundamental ring of $p$-adic Hodge Theory $A_{inf}$ (introduced by Jean-Marc Fontaine in (Fontaine, 1994)).
@@ 1.4 Acknowledgments
I began thinking about the problem addressed in this paper while I was on a sabbatical (Spring 2018) and enjoying the support and hospitality of Research Institute of Mathematical Sciences (RIMS), Kyoto and I thank RIMS for the same. I thank Shinichi Mochizuki for conversations around many topics of common interest during my visit. I thank Yuichiro Hoshi for conversations about Anabelian Reconstruction Theory. I thank Machiel van Frankenhuijsen for many conversations related to Mochizuki’s work and the $abc$-conjecture. I thank Taylor Dupuy and Anton Hilado for some conversations surrounding Mochizuki’s works and for pointing out some typos in an earlier version of this paper. Finally it is a pleasure to thank Michael Hopkins for his insightful lectures at the Arizona Winter School 2019 and the organizers of the School for an excellently organized meeting.
@@ 2.3 Existence of formal groups with monoid actions
Let ${\mathbb{Q}}[M]$ be the monoid ring of a monoid $M$ over ${\mathbb{Q}}$. Let $R\supset{\mathbb{Q}}[M]$ be a commutative ring. Then I claim that there always exists an $M$-formal group over $R$.
@@ 4.2 Perfectoid rings, fields and Huber pairs and monoid formal group laws
This demonstrates the close relationship between Mochizuki’s idea of variation of ring structures keeping multiplicative structures fixed and the ideas in perfectoid geometry of (Scholze, 2012), (Kedlaya and Liu, 2015) (the relationship with (Fargues and Fontaine, 2018) is detailed in (Joshi, 2023a) and will not be discussed here). This recognition played an important role in the theory of Arithmetic Teichmuller Spaces documented extensively in (Joshi, 2021, 2023b, 2023a, 2024b). In this section, I record these observations.
@@
4.3 Global arithmetic analog of Fontaine’s $p$-adic ring $A_{inf}$
Let ${\mathbb{C}}_{p}^{\flat}$ be the tilt of ${\mathbb{C}}_{p}$ ((Scholze, 2012)). In the perfectoid setting, and modern $p$-adic Hodge Theory ((Fargues and Fontaine, 2018), (Scholze, 2012), (Kedlaya and Liu, 2015) and others) the ring $A_{inf}=W({\mathscr{O}}_{{\mathbb{C}}_{p}^{\flat}})$ introduced in (Fontaine, 1994) plays a fundamental role. I want to remark that the results of Sections 4.1, 4.2 especially Theorem 4.1.1, Theorem 4.2.1 and Corollary 4.2.2 suggest a definition of a global i.e. a number field analog
$\mathbb{A}_{inf,L}$
of Fontaine’s ring ((Fontaine, 1994))
$A_{inf}=W({\mathbb{C}}_{p}^{\flat})$
in the $p$-adic setting and a definition of untilts of a number field. Let me provide a brief sketch of the construction of this global ring (detailed study of the properties of $\mathbb{A}_{inf,L}$ and its relation with Mochzuki’s global arithmetic Kodaira-Spencer classes ((Mochizuki, 1999)) is taken up in (Joshi, 2024a) which is under preparation).
For a ring $R$, let $\mathbb{W}(R)$ be the ring of big Witt-vectors of $R$ ((Hazewinkel, 1978)); let $L$ be a number field, let $\bar{L}$ be an algebraic closure, let ${\mathscr{O}}_{L}$ (resp. ${\mathscr{O}}_{\bar{L}}$) be the ring of integers of $L$ (resp. $\bar{L}$). Let
[…]
be the perfection of the multiplicative monoid of the ring of integers ${\mathscr{O}}_{\bar{L}}$ of $\bar{L}$. This is the notion of perfection of monoids introduced in (Mochizuki, 2008) and hence this means $M$ is a perfect multiplicative monoid in the sense of (Mochizuki, 2008).
Note that one has a natural action $G_{L}\curvearrowright M$ of the absolute Galois group $G_{L}$ on $M$, as well as surjective Frobenius morphisms $x\longmapsto x^{n}$ for all $n\geq 1$. An element of $M$ can described as a compatible system of $n^{th}$-roots (all $n\geq 1$) of some $x\in{\mathscr{O}}_{\bar{L}}$.
Let ${\mathbb{Z}}[M]$ be the monoid ring of $M$. Then Theorem 4.1.1, Theorem 4.2.1 and Corollary 4.2.2 suggest that
(4.3.2) $\mathbb{A}_{inf,L}=\mathbb{W}_{L}=\mathbb{W}({\mathbb{Z}}[M])\otimes_{{\mathbb{Z}}}{\mathscr{O}}_{L}=\mathbb{W}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])\otimes_{{\mathbb{Z}}}{\mathscr{O}}_{L}$
equipped with
  1. (1)
    a surjective homomorphism of rings
    $\mathbb{W}_{L}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])\to{\mathscr{O}}_{\bar{L}},$
    (so the kernel is a prime ideal $P=\ker(\mathbb{W}_{L}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])\to{\mathscr{O}}_{\bar{L}})$),
  2. (2)
    the given structure (4.3.2) of ${\mathscr{O}}_{L}$-algebra on $\mathbb{W}_{L}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])$,
  3. (3)
    the Teichmuller-representative homomorphism of a perfect multiplicative monoid
    $M={\mathscr{O}}_{\bar{L}}^{pf}\to\mathbb{W}_{L}({\mathbb{Z}}[M])=\mathbb{W}_{L}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])$
    such that the composite homomorphism of multiplicative monoids
    $M={\mathscr{O}}_{\bar{L}}^{pf}\to\mathbb{W}_{L}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])\to{\mathscr{O}}_{\bar{L}}$
    induces an isomorphism on the passage to perfections of monoids;
  4. (4)
    Frobenius endomorphisms induced from the Frobenius endormorphisms $x\longmapsto x^{n}$ (for each $n\geq 1$) of $\mathbb{W}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}])$;
  5. (5)
    the canonical action of the absolute Galois group $G_{L}$ (through its action on $M$),
is the global analog of $A_{inf}$. The set of “untilts of $\bar{L}$” should be a Galois, Frobenius stable subset of ${\rm Spec}(\mathbb{W}_{L}({\mathbb{Z}}[{\mathscr{O}}_{\bar{L}}^{pf}]))$ modulo the action of ${\mathscr{O}}_{L}^{*}$ (by (4.3.2)). This ring is studied in (Joshi, 2024a).
The following assertion is now clear:
Theorem 4.3.3.
Let $L$ be a number field, $\bar{L}$ and an algebraic closure of $L$, $M$ be the perfection of the multiplicative monoid ${\mathscr{O}}_{\bar{L}}$ defined in eq. (4.3.1) and let the other notations of this subsection be in force. Then there is a natural homomorphism
$L_{M}\to\mathbb{A}_{inf,L},$
from the universal ring $L_{M}$ provided by the universal formal group law of Theorem 4.1.1 for the monoid $M$, to the ring $\mathbb{A}_{inf,L}$ defined by eq. (4.3.2).
Proof.
The proof is clear from the proofs of Theorem 4.1.1 and Theorem 4.2.1. ∎